OpenStudy (anonymous):

How many moles of sucrose are there in 2.50L of a 0.500 molar solution?

OpenStudy (accessdenied):

The definition of molarity is moles per liter. If we multiply the volume and molarity, we will get the moles. \( \text{moles} = \dfrac{\text{moles}}{\text{liters}} \times \text{liters} = M \times V \)